OpenStudy (anonymous):

Simplify sin X (1 + cot2 X ).

OpenStudy (anonymous):

\[\sin x \left( 1+\cot 2x \right)=\sin x \left( 1+\frac{ \cos 2x }{\sin 2x } \right)\] \[=\sin x \frac{ \sin 2x+\cos 2x }{\sin 2x }=\sin x \frac{ 2\sin x \cos x+\cos ^{2}x-\sin ^{2} x }{2\sin x \cos x }\] \[=\frac{ 2\sin x \cos x+\cos ^{2}x-\sin ^{2}x }{2\cos x }\]

OpenStudy (raden):

is that cot(2x) or cot^2 (x) ?

OpenStudy (anonymous):

X= theta